# AdaptiveDrag: Semantic-Driven Dragging on Diffusion-Based Image Editing

### v1

@@ 1 Introduction
  • We prose Correspondence Sample to improve the generation stability of the diffusion process, encouraging the semantic consistency between regions of handle and target points.
@@ 2.2 Diffusion-Based Image Editing
Recently, the impressive generative capabilities of large-scale text-to-image diffusion models have led to the development of numerous methods based on these models Rombach et al. (2022); Saharia et al. (2022). For interactive image editing, DragDiffusion Shi et al. (2024) employs a point-based image editing scheme based on the diffusion model, similar to DragGAN. This method utilizes LoRA for identity-preserving fine-tuning and optimizes the latent space using the loss function of motion supervision and point tracking. However, as shown in Fig. 1, previous methods (e.g., DragDiffusion, EasyDrag) face two main issues: ‘drag missing’ and ‘feature maintenance failure’ which result in the latent being incorrectly positioned in certain regions. FreeDrag Ling et al. (2024) introduces a template feature through adaptive updating and line search with backtracking strategies, resulting in more stable dragging. DragNoise Liu et al. (2024) presents a semantic editor that modifies the diffusion latent in a single denoising step, leveraging the inherent bottleneck features of U-Net. Nevertheless, these methods still have challenges when dragging over long distances or across complex textures. To design a user-friendly point-based image editing method, EasyDrag Hou et al. (2024) leverages gradients in the motion supervision process that remain unchanged in areas with small gradients, and it automatically generates the mask M. However, gradient information is not always directly related to the image content, which can result in inaccurate mask generation and unsatisfactory image outcomes. In contrast to previous work, we propose a novel semantic-driven point-based image editing framework that achieves precise results across different drag ranges without the need for a mask.
@@ 4.1 Implementation Details
DragDiffusion  Shi et al. (2024) DragNoise  Liu et al. (2024) EasyDrag  Hou et al. (2024) AdaptiveDrag (Ours)
Conference CVPR 2024 CVPR 2024 CVPR 2024 -
MD $\downarrow$ 34.29 40.89 34.44 30.69
IF (1-LPIPS) $\uparrow$ 0.789 0.861 0.882 0.873

### v2

@@ 1 Introduction
  • We propose Correspondence Sample to improve the generation stability of the diffusion process, encouraging the semantic consistency between regions of handle and target points.
@@ 2.2 Diffusion-Based Image Editing
Recently, the impressive generative capabilities of large-scale text-to-image diffusion models have led to the development of numerous methods based on these models Rombach et al. (2022); Saharia et al. (2022). For interactive image editing, DragDiffusion Shi et al. (2024) employs a point-based image editing scheme based on the diffusion model, similar to DragGAN. This method utilizes LoRA for identity-preserving fine-tuning and optimizes the latent space using the loss function of motion supervision and point tracking. However, as shown in Fig. 1, previous methods (e.g., DragDiffusion, EasyDrag) face two main issues: ‘drag missing’ and ‘feature maintenance failure’ which result in the latent being incorrectly positioned in certain regions. FreeDrag Ling et al. (2024) introduces a template feature through adaptive updating and line search with backtracking strategies, resulting in more stable dragging. DragNoise Liu et al. (2024) presents a semantic editor that modifies the diffusion latent in a single denoising step, leveraging the inherent bottleneck features of U-Net. Nevertheless, these methods still have challenges when dragging over long distances or across complex textures. To design a user-friendly point-based image editing method, EasyDrag Hou et al. (2024) leverages gradients in the motion supervision process that remain unchanged in areas with small gradients, and it automatically generates the mask M. Moreover, some methods shi2024instadrag; shin2024instantdrag; lu2024regiondrag; cui2024stabledrag attempt to improve the quality of results in various ways (e.g., drag by regions lu2024regiondrag, flow-based drag shin2024instantdrag, and the fast editing method shi2024instadrag). However, these previous methods provided the mask is not always directly related to the image content, which can result in inaccurate mask generation and unsatisfactory image outcomes. In contrast to previous work, we propose a novel semantic-driven point-based image editing framework that achieves precise results across different drag ranges without the need for a mask.
@@ 4.1 Implementation Details
DragDiffusion  Shi et al. (2024) DragNoise  Liu et al. (2024) EasyDrag  Hou et al. (2024) FastDrag  Zhao et al. (2024) AdaptiveDrag (Ours)
Conference CVPR 2024 CVPR 2024 CVPR 2024 NeruIPS 2024 -
MD $\downarrow$ 34.29 40.89 34.44 34.13 30.69
IF (1-LPIPS) $\uparrow$ 0.789 0.861 0.882 0.859 0.873
@@ A.4 Reproducibility Statement
Figure 15: A failure case of our approach. We attempt to drag the boy’s shoulder to a higher position, while the entire body becomes unexpectedly expanded.
A.5 Time Consumption
In this section, we present the time consumption in Tab. 3. Compared with the manual mask, our auto mask network only requires 3.2s for generating a mask region. The time cost of manual methods is about 16.6s, which is obtained from the average of ten images edited by each of the five users. These results verify the efficiency of our auto mask design.
Table 3: Time consumption of DragDiffusion and AdaptiveDrag using images from the DragBench Shi et al. (2024) dataset. The experiment is conducted on a single Nvidia V100 GPU, with input images size are 512 $\times$ 512.
Method Mask LoRA Optimization Sample
AdaptiveDrag SAM SLIC 40.1s 10.3s 20.4s
3.0s 0.2s
DragDiffusion 24.9s 39.9s 31.2s 5.1s
Shi et al. (2024)
A.6 Limitations
Fig. 15 illustrates the failure case of our method. AdaptiveDrag has limitations in editing the image content in ways that do not align with real-world scenarios (e.g., moving only the boy’s shoulder to higher positions). This is primarily due to the pre-trained diffusion model, which incorporates basic rules that are consistent with real-world scenes.
A.7 Stability Analysis
In this section, we adopt various operations in the same scene. As shown in Fig. 16, the AdaptiveDrag method achieves stable and high-quality results with expand, move, and resize operations, demonstrating the robustness and stability of our approach.
Figure 16: Various operations in the same scene (expand, moving, resize) based on the DragBench Shi et al. (2024) dataset. Our method has stability with a different point-based editing.
A.8 More Details of AdaptiveDrag
To facilitate a better understanding, we provide pseudocode for Section 3.4.2 as follows. The whole pipeline if AdaptiveDrag is present in Alg. 1. Following DragDiffusion Shi et al. (2024), the process of the motion supervision and point tracking are provided in Eq. 5 and Eq. 6.
Input: Input image $z_{0}$, handle point $\{p_{i}^{0}\}_{i=1}^{l}$, target point $\{q_{i}\}_{i=1}^{l}$, drag iterations $N$, latent time steps $T$, number of segmentation patches for SLIC ${n_{p}}$
Output: Output image $z_{0}^{\prime}$
1 Finetune $U_{l}$ on $z_{0}$ with LoRA
2Generate mask $M$ with SAM and SLIC (Sec. 3.3)
3Superpixel segmentation patches $\{A_{(x_{i},y_{i})}\}_{i=1}^{l}$
[…]
6for $k$ in 0:K-1 do
7       $z_{T,0}^{k}\leftarrow z_{T}^{k}$
8      Update $z_{i}^{k}$ using motion supervision with patch $A_{(x_{i},y_{i})}$ as Eq. 5
9      Update $p_{i}^{k+1}$ using points tracking with patch $A_{(x_{i},y_{i})}$ as Eq. 6
10      Calculate the updating distance $d^{u}_{i}\leftarrow$ $p_{i}^{k+1}-p_{i}^{k}$
11      Ideal Distance $d^{k}_{i}\leftarrow\frac{D_{i}}{N}$ (D is the distance from $p_{i}^{0}$ to $q_{i}$)
12      if $d^{u}_{i}<{d^{k}_{i}}$ then
13             $z_{i}^{k}\leftarrow z_{i}^{k-1}$
14            $p_{i}^{k+1}\leftarrow p_{i}^{k}$
15       end if
16      for t in T:1 do
17             $z_{t-1}^{K}\leftarrow$ denoising from $z_{t}^{N}$ (Eq. 2)
18       end for
19      $z^{\prime}_{0}\leftarrow z_{0}^{K}$
20 end for
Algorithm 1 Pipeline of AdaptiveDrag
The motion supervision process of the latent $z_{t}^{k}$ can be formulated as:
$\mathcal{L}_{\mathrm{ms}}(z_{t}^{k})$ $=\sum_{i=1}^{l}\sum_{q\in A(x_{i},y_{i})}\left\|F_{q+d_{i}}(z_{t}^{k})-\mathrm{sg}(F_{q}(z_{t}^{k}))\right\|_{1}+\lambda\left\|(z_{t}^{k-1}-\mathrm{sg}(z_{t}^{k-1}))\odot(1-M)\right\|_{1},$ (5)
where $z_{t}^{k}$ is the $t$-th step latent after $k$-th step optimization, $sg(\cdot)$ is the stop gradient operator and $M$ is the mask region from auto mask generation network. The $F(z)$ is the output feature of the Diffusion UNet. We denote the superpixel patch centered around $p_{i}^{k}$ as $A(x_{i},y_{i})$,
The update process of handle points can be formulated as:
$p_{i}^{k+1}=\underset{q\in A(x_{i},y_{i})}{\operatorname*{\arg\min}}\left\|F_{q}({z}_{t}^{k+1})-F_{p_{i}^{0}}(z_{t})\right\|_{1}.$ (6)
A.9 Additional Methods Comparison
In this section, we present additional comparisons between our AdaptiveDrag and other state-of-the-art methods, as illustrated in Fig. 17.
Figure 17: Additional compared methods (FastDrag, InstantDrag, Region Drag) based on the DragBench Shi et al. (2024) dataset. Our method also delivers more precision and high-quality results.
A.10 More Analysis of Correspondence Sample
In this section, we present the visual comparison of the “Correspondence Sample” design. As shown in Fig. 20, compared with the method without “Correspondence Sample”, our AdaptiveDrag generates better quality results, which contain content more aligned with the target point locations.
Figure 18: The visual comparison of the correspondence sample design. Specifically, ’w/o CLoss’ denotes the method without CLoss in the sample stage, while ’w/ CLoss’ represents our approach. AdaptiveDrag with CLoss optimization effectively edits the mountains into the desired target locations, whereas the method without CLoss fails to achieve this.
A.11 Additional Figures about Mask Generation
Figure 19: The visual comparison of the generated mask area with different editing points.
Figure 20: The visual results of the generated mask with the network only contain the SAM or the SLIC model.